Steam at 100 bar, 3200 C expands isothermally in a cylinder behind a piston to a pressure of 10 bar. Solve for the work done per kilogram of steam. 


1
Expert's answer
2021-03-03T11:24:04-0500

"W=\frac{m}{M}RT\ln\frac{p_1}{p_2}\to W/m=\frac{RT}{M}\ln\frac{p_1}{p_2}="


"=\frac{8.31\cdot 593}{0.018}\ln\frac{100}{10}=630375\ (J/kg)" . Answer